OpenStudy (anonymous):

PLEASE HELP! I am not sure how to use the vertical line test on x-y^2=0

OpenStudy (***[isuru]***):

make y subject... y^2 = x y = plus or minus sqrt(x) \[y= \pm \sqrt{x}\] the graph will be like http://fooplot.com/plot/rpvdqg70cs the u can try the vertical line test.... lol hope this will help ya!!

OpenStudy (anonymous):

lol I have the graph. I am not sure how to do the vertical line test and whether y is a function of x.

OpenStudy (anonymous):

I think I figured that one out. Y is not a function of x, right?

OpenStudy (anonymous):

Could you help me on this one? x^2=xy-1 is y a function of x?

OpenStudy (***[isuru]***):

yes... vertical line test says that if y is a function of a x the the vertical line cannot be cut more than once when u r moving it along the x axis... here it's get cut twice i mean for ur 1st one

OpenStudy (***[isuru]***):

again make y the subject x^2 = xy -1 y = ( x^2 - 1)/x where x is not equal to 0 u can do the vertical line test here.. it's a function of x

OpenStudy (***[isuru]***):

though it seems a little confusing when x becomes closer to 0 they never intersect so ...it's a function of x http://fooplot.com/plot/j4f076ghtt